$SPAD/src/input dfloat.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool dfloat.output
)set message test on
)set message auto off
)clear all
--S 1 of 10
machineFraction(2.71828)
--R
--R
--R        382563899846521
--R   (1)  ---------------
--R        140737488355328
--R                                                      Type: Fraction(Integer)
--E 1
--S 2 of 10
machineFraction(2.71828@DoubleFloat)
--R
--R
--R        382563899846521
--R   (2)  ---------------
--R        140737488355328
--R                                                      Type: Fraction(Integer)
--E 2
--S 3 of 10
machineFraction(2.71828 :: DoubleFloat)
--R
--R
--R        382563899846521
--R   (3)  ---------------
--R        140737488355328
--R                                                      Type: Fraction(Integer)
--E 3
--S 4 of 10
machineFraction(eApprox : DoubleFloat := 2.71828)
--R
--R

--R        382563899846521
--R   (4)  ---------------
--R        140737488355328
--R                                                      Type: Fraction(Integer)
--E 4
--S 5 of 10
avg : List DoubleFloat -> DoubleFloat
--R
--R                                                                   Type: Void
--E 5
--S 6 of 10
avg l ==
  empty? l => 0 :: DoubleFloat
  reduce(_+,l) / #l
--R
--R                                                                   Type: Void
--E 6
--S 7 of 10
avg []
--R
--R   Compiling function avg with type List(DoubleFloat) -> DoubleFloat
--R
--R   (7)  0.
--R                                                            Type: DoubleFloat
--E 7
--S 8 of 10
machineFraction(avg [3.4,9.7,-6.8])
--R
--R
--R        1182194902184755
--R   (8)  ----------------
--R         562949953421312
--R                                                      Type: Fraction(Integer)
--E 8
--S 9 of 10
machineFraction(cos(3.1415926)$DoubleFloat)
--R
--R
--R          9007199254740979
--R   (9)  - ----------------
--R          9007199254740992
--R                                                      Type: Fraction(Integer)
--E 9
--S 10 of 10

machineFraction(cos(3.1415926 :: DoubleFloat))
--R
--R
--R           9007199254740979
--R   (10)  - ----------------
--R           9007199254740992
--R                                                      Type: Fraction(Integer)
--E 10
)spool
)lisp (bye)

———-

References^{[1] nothing}

6